OpenStudy (anonymous):

Factor the trinomial. x^2 + 4x - 32

OpenStudy (hoblos):

x^2 + 4x - 32 =x^2 +8x -4x -32 = x(x+8) - 4(x+8) = (x+8)(x-4)

OpenStudy (yrivers36):



OpenStudy (yrivers36):

sorry so sloppy